OpenStudy (anonymous):

During a hike, Vijay saw a total of 19 birds and butterflies. Knowing that birds have 2 wings and butterflies have 4 wings, he counted 52 wings. Use a model to determine how many birds he saw. 9 10 11 12

OpenStudy (anonymous):

12

OpenStudy (zzr0ck3r):

a = birds b = butterflys a+b=19 2a+4b=52 so a=19-b 2(a)+4b=52 2(19-b)+4b=52 38-2b+4b=52 2b=14 b=7 so a = 12

OpenStudy (anonymous):

Thanks